OpenStudy (anonymous):

Laplace transform: L{e^(2t) * sin(10t)}?

OpenStudy (anonymous):

i would help you if i knew how to. :l sorry!

OpenStudy (amistre64):

use the definition

OpenStudy (amistre64):

\[\int e^{-st}e^{2t}sin(10t)dt\to\ \int e^{-(s-2)t}sin(10t)dt\]

OpenStudy (amistre64):

u = s-2 \[\int e^{-ut}sin(10)dt=\frac{10}{u^2+10^2}\to\ \frac{10}{(s-2)^2+10^2}\]

OpenStudy (anonymous):

Thank you!

OpenStudy (amistre64):

yw

OpenStudy (amistre64):

with any luck i remembered the tabling for the sin correctly :)

OpenStudy (anonymous):

I didn't remember it was one of the basic transforms :P